Archive for August, 2006

Generator Installation

Posted: Wednesday, August 30th, 2006 at 17:14 by Neil Watson

Given recent power issues we have been sourcing new capacity in relation to our backup generator supply. We are pleased to inform you that we have been advised that a single new 900kVA generator will be in place for Friday morning, when it can then be connected, without interruption, to our supply.We are currently planning for the work to begin at 08.00hrs on Friday morning; the work can be carried out whilst the main power supply is connected, therefore providing us with the opportunity to complete without downtime.

There is however a small risk associated with this method, as during the work we will be without back up generator. The work is envisaged to take less than 30 minutes to complete. Following on from the installation, the generator will be fully tested on Sunday morning, we are currently waiting to confirm the time of the test and will advise as soon as this is available.

In the near future we will be adding further capacity to our back up power and will add details to noc.enta.net as appropriate.

  1. Neil Watson says:

    There has unfortunately been a delay in the work being carried out as described above. We are currently waiting for a confirmed time for the switchover and will advise as soon as possible.

  2. Neil Watson says:

    We have now confirmed that a full load test of the generator will take place this coming Sunday, 03rd Deptember 2006 at 08.00hrs.

ADSL Speed Issues

Posted: Wednesday, August 30th, 2006 at 12:23 by Neil Watson

Some ADSL end users are currently experiencing speed issues when connected to a selection of our LNS routers. We are currently investigating the issue and will report further information as soon as possible.

  1. Neil Watson says:

    ADSL speeds have now returned to normal for all customers. The problems were caused by a misconfigured LNS router that was announcing several routes, causing a multitude of traffic to pass through a router not designed for this load. The routes in question have now been removed from the router and service has been returned to normal. We again apologise for any inconvenience that the slow speeds may have caused.

Framestream router upgrade

Posted: Wednesday, August 30th, 2006 at 09:30 by Neil Watson

Date: 05th September 2006 00.00 – 02.00hrs

As part of our ongoing network monitoring, maintenance and improvements, we will be replacing routing equipment in Telehouse East this coming Monday night. The work involved is minimal, as the majority of tasks will be performed whilst the replacement equipment is offline; however a small period of downtime, in the region of a couple of minutes, will be required for all framestream customers. The swap over to the new 7206 NPE G1 router will be performed between midnight at 02.00hrs, during which Framestream customer should expect to see an a small outage on their circuit. We apologise for any inconvenience this may cause and wish to assure you that all steps are being taken to keep any disruption to a minimum.

IOS Upgrade- Router reload

Posted: Tuesday, August 22nd, 2006 at 13:01 by Neil Watson

We will be reloading the following router for an emergency software update.telehouse-east2.bdr.enta.net at approx 04:00hrs on 23rd August 2006
The expected downtime for the reload is 5 mins, during which, customers who are directly connected to this unit will suffer a small period of downtime. Other customers should not be affected, although there is a chance of some routing instability for a few minutes after the reload. We apologise for any inconvenience this may cause.

  1. Steve Lalonde says:

    this was sent out with the wrong router name originaly it is telehouse-east2.bdr.enta.net that will be rebooted and NOT telehouse-east.bdr.enta.net.

  2. Neil Watson says:

    Problems were unfortunately experienced during this work, causing a greater period of downtime than originally envisaged. Whilst the majority of customers will not have been affected, anyone directly connected to the router will have been without service until approximately 07.30 this morning. We currently have engineers on site in Telehouse who are monitoring the situation. At this present moment, we do not expect any further downtime from this work. We again apologise to the small percentage of customers affected.

Problems contacting Support

Posted: Friday, August 18th, 2006 at 14:29 by Neil Watson

We are unfortuantley experiencing some issues with out main support number, 0870 224 3494, resulting in customers not being connected to our support team. Our team of engineers are currently looking into this issue for us. We apologise for any inconvenience caused and would advise you to email your request if possible. Updates will be made as they become available.

  1. Neil Watson says:

    All phone issues have now been resolved. Our support team is once again available on 0870 224 3494. We again apologise for any invonience.

LINX port upgrade

Posted: Thursday, August 17th, 2006 at 14:13 by Neil Watson

Date: 18th August 2006, 03:00 – 05:00hrs

We will be upgrading one of our two 1Gb LINX ports to 10Gb during the above period. The existing port will be shut down and all traffic will reroute via the other port. There should be no noticeable effect during the work and no outage is envisaged.

Cambridge POP

Posted: Thursday, August 17th, 2006 at 13:51 by Neil Watson

We are currently aware of power issues affecting Telehouse North, whilst there is no direct link between the affected areas and Entanet, there is a risk that one of our suppliers may experience problems. As such, we would like to inform all customers that connect to our Cambridge POP that there is a danger of losing connectivity. The links to this POP are protected circuits; therefore, the threat is low, however, the risk does remain. We will update as soon as more information is available.

Incident report: Power outage 16th August 2006

Posted: Thursday, August 17th, 2006 at 13:12 by Neil Watson

Please find the below incident report regarding the power issues experience at out Telford HQ on 16th August 2006. We apologise for any inconvenience this caused and wish to reassure you that we are taking all necessary action to prevent any reoccurrence.

Incident report – Power Outage 16.08.2006

E-mail delays

Posted: Friday, August 11th, 2006 at 12:25 by Neil Watson

Our dedicated systems team are aware of a current mail issue that will be affecting a proportion of mail that passes through our systems. Due to the amount of mail processed daily, we have steps in place that prevent any of our mail servers from becoming overwhelmed, this is an automated process that stops individual servers in the mail server farm from accepting SMTP mail until queues have processed. We currently have one server, where the ‘blocking script’ ceased to function; hence, the mail server currently has a large backlog of mail to process. The script is once again running, but time is needed for the backlog to dissipate. This may cause some delays to mail being processed through this server. We apologise for any inconvenience caused.

Upgrade of LNS Server IOS

Posted: Wednesday, August 9th, 2006 at 12:34 by Neil Watson

Planned: 10th August 2006, 03.15hrs

Following on from recent authentication issues, it is necessary to upgrade the IOS (Operating System) on one of our LNS servers. This work will require a reboot of the server, which will unfortunately disconnect any user on the affected server at the time. We expect this to be a small percentage of our total users and all connections should reconnect as a matter of course to other available servers. This work is scheduled out of hours to cause the minimum amount of disruption to our customers.